A motor car of mass 800kg travelling at 20ms(on top of it there is -1)is brought to rest by brakes in 100m.
Calculate the average braking force required
1
Expert's answer
2020-08-28T09:33:25-0400

Let us write equations of motion of the car, assuming it moves with constant retardation (and hence constant braking force):

"S(t) = v_0 t - \frac{a t^2}{2}", "v(t) = v_0 - a t".

At the moment of time, when the car stops completely "v(t') = v_0 - a t' = 0", from where the time it takes to completely stop is "t' = \frac{v_0}{a}".

The horizontal distance covered in time "t'" is given ("L = 800m)", therefore: "L = S(t') = \frac{v_0^2}{a} - \frac{v_0^2}{2 a} = \frac{v_0^2}{2 a}", from where the acceleration is "a = \frac{v_0^2}{2 L}", and from second Newton's law "a = \frac{F}{m}", so "\frac{v_0^2}{2 L} = \frac{F}{m}", from where the force is "F = \frac{m v_0^2}{2 L} = 1600 N".